OpenStudy (anonymous):

6. Iodine chloride, ICl, can be made by the following reaction between iodine, I2, potassium iodate, KIO3, and hydrochloric acid. 2 I2 + KIO3 + 6 HCl ---------> 5 ICl + KCl + 3 H2O Calculate how many grams of iodine are needed to prepare 28.6 grams of ICl by this reaction.

OpenStudy (jfraser):

a balanced reaction gives us the mole ratio of pieces that will react. You need to use the molar mass of ICl and the 28.6g to find the moles of ICl you want to make. That moles of product will be made by a molar ratio of iodine, I2. That many moles of iodine will have a mass proportional to the molar mass of iodine\[28.6g ICl * (\frac{1mol ICl}{162.4g ICl})*(\frac{2mol I_{2}}{5mol ICl})*(\frac{253.8g I_{2}}{1mol I_{2}})\] I get about 17.8g of I2.